Say true or false and justify your answer: 
10 × 1011 = 10011

Grade:12

1 Answers

Harshit Singh
2 years ago
Dear Student

Let us consider Left Hand Side (LHS) = 10 × 10^11
10^1 + 11= 10^12 ... [∵a^m × a^n = a^m + n]
Now, consider Right Hand Side (RHS) = 100^11
= (10 × 10)^11
= (10^1 + 1)^11
= (10^2)^11
= (10)^2x11 ...[∵(a^m )^n = a^mn ]
=(10)^22
Thanks